# Blackwell’s Approachability with Approximation Algorithms

### v1

@@ 1 Introduction
Algorithms for choosing the player’s actions so that approachability to the target set is guaranteed (when possible), naturally require, as a sub-procedure, to solve certain optimization problems with respect to the actions sets of the player and the adversary (and also the target set $S$). In this work however, we are interested in the case in which such optimizations may be computationally-infeasible. For instance, when one of these sets, $\mathcal{X}$ or $\mathcal{Y}$ (or both), corresponds to the set of all possible solutions to some NP-Hard optimization problem (or more accurately, the convex hull of all such solutions, to include mixed-actions). As an example, we can think that the action the player needs to take on each round of the game is to select a vertex-cover or a Hamiltonian cycle in some fixed combinatorial graph, or that the adversary’s action corresponds to choosing a cut in some fixed combinatorial graph. In such a case, while the Blackwell instance might be approachable, it might not be efficiently approachable, since no polynomial-time algorithm that can guarantee approachability is known (or even widely expected to exist). Instead, in such a case it is natural to assume that the action sets (those for which optimization is difficult) are accessible through a computationally-efficient approximation algorithm (approximation oracle henceforth) with some ratio $\alpha>0$. If $\alpha>1$, then for any linear cost function, the oracle returns a point in the action set whose value is at most $\alpha$ times that of the action with minimal value; if $\alpha<1$, the oracle returns a point in the action set whose value is at least $\alpha$ times that of the action with maximal value. The question is then, what can the player guarantee (w.r.t. his average loss) using only a polynomial (in the natural parameters of the problem) number of calls (and additional polynomial running time) to these oracles.
@@ 6 Approximation Oracle for the set $\mathcal{Y}$
Using again the bi-linearity of $g$ is bi-linear, we thus have that

### v2

@@ 1 Introduction
Algorithms for choosing the player’s actions so that approachability to the target set is guaranteed (when possible), naturally require, as a sub-procedure, to solve certain optimization problems with respect to the actions sets of the player and the adversary (and also the target set $S$). In this work however, we are interested in the case in which such optimizations may be computationally-infeasible. For instance, when one of these sets, $\mathcal{X}$ or $\mathcal{Y}$ (or both), corresponds to the set of all possible solutions to some NP-Hard optimization problem (or more accurately, the convex hull of all such solutions, to include mixed-actions). As an example, we can think that the action the player needs to take on each round of the game is to select a vertex cover or a Hamiltonian cycle in some fixed combinatorial graph, or that the adversary’s action corresponds to choosing a cut in some fixed combinatorial graph (see a more detailed example in the sequel). In such a case, while the Blackwell instance might be approachable, it might not be efficiently approachable, since no polynomial-time algorithm that can guarantee approachability is known (or even widely expected to exist). Instead, in such a case it is natural to assume that the action sets (those for which optimization is difficult) are accessible through a computationally-efficient approximation algorithm (approximation oracle henceforth) with some ratio $\alpha>0$. If $\alpha>1$, then for any non-negative linear objective, the oracle returns a point in the action set whose value is at most $\alpha$ times that of the action with minimal value; if $\alpha<1$, the oracle returns a point in the action set whose value is at least $\alpha$ times that of the action with maximal value. The question is then, what can the player guarantee (w.r.t. his average loss) using only a polynomial (in the natural parameters of the problem) number of calls to these oracles (and additional polynomial running time).
As one illustrative example, we can consider the following game in which there is a fixed undirected graph $G(V,E)$, where the set of vertices $V$ is partitioned into two subsets $V_{1},V_{2}$, each is controlled by a different interested party. The player, representing both parties, is required on each iteration to choose action $x\in\mathcal{X}\subset[0,1]^{|V|}$ which is a distribution (to allow mixed actions) over vertex covers of the graph333here a vertex cover is represented by its identfiyng vector which is an element in $\{0,1\}^{|V|}$. The adversary in turn chooses $y\in\mathcal{Y}\subset\mathbb{R}^{|V|}_{+}$ which assigns a (bounded) non-negative weight to each vertex. The bilinear loss is then a vector in +$\mathbb{R}^{2}_{+}$ which assigns to each party $i\in\{1,2\}$ its weighted loss, i.e., $\ell(x,y)=(\sum_{j\in{}V_{1}}x_{j}y_{j},\sum_{j\in{}V_{2}}x_{j}y_{j})^{\top}$. The pre-defined target set $S$ is designed to balance between the interests of both parties. While minimum-weight vertex cover is NP-Hard, there are well known efficient $2$-approximation algorithms, e.g., [2].
@@ 6 Approximation Oracle for the set $\mathcal{Y}$
Using again the bi-linearity of $g$, we thus have that
@@ 7.2 Putting everything together
Acknowledgements
Dan Garber is supported by the ISRAEL SCIENCE FOUNDATION (grant No. 2267/22).